# When Do Tools and Planning Help LLMs Think? A Cost- and Latency-Aware Benchmark

### v1

@@ Abstract
Modern large language models (LLMs) increasingly rely on inference-time planning and external tools to improve reasoning. We benchmark this behavior on two real-world settings: event-centric question answering over graph-structured knowledge (Event-QA) and persuasive response generation in Reddit ChangeMyView (CMV). Using LangChain and LangGraph, we compare a one-shot baseline against a plan–execute–replan agent equipped with task-specific tools (DBpedia SPARQL/lookup/schema exploration, Wikipedia-focused retrieval, and topical web search). We evaluate on 60 examples each from Event-QA and CMV (3 splits of 20), and report both mean end-to-end latency and per-example token cost estimates. We evaluate GPT-4o and GPT-4o-mini under identical workflows and report accuracy and end-to-end latency. On Event-QA, the best tool-augmented configuration improves accuracy (e.g., 47.5% $\rightarrow$ 67.5% for GPT‑4o) while increasing latency by orders of magnitude ($\sim$8s $\rightarrow$ $\sim$317s per example). On CMV, one-shot prompting is strongest (e.g., GPT‑4o‑mini achieves 75% at $\sim$6s), and planning+search increases latency substantially without consistent gains. However, complex multi-tool orchestration exposes failure modes where the smaller model degrades. Overall, the findings highlight the need for task-specific, cost-aware choices of both model size and agent/tooling complexity.
@@ I Introduction
Large Language Models (LLMs) based on Transformer architectures have rapidly evolved into general-purpose systems for text understanding, generation, and reasoning [vaswani2017attention, wang2024historyllm]. Beyond surface-level language fluency, recent development has emphasized the ability of LLMs to perform multi-step reasoning (planning intermediate steps and integrating external evidence)—i.e., to perform multi-step reasoning, plan intermediate steps, and reliably integrate external information sources when answering complex real-world questions. A common and influential approach is Chain-of-Thought (CoT) prompting, which elicits intermediate reasoning traces that can substantially improve performance on multi-step problems [wei2022chain_of_thought]. Subsequent work has shown that inference-time strategies such as sampling multiple reasoning paths and aggregating consistent answers (self-consistency) can further improve reliability without changing model parameters [wang2022self_consistency]. More broadly, recent lines of research argue that test-time compute—allocating additional deliberation or search at inference—can sometimes yield improvements comparable to or greater than scaling model size [snell2024test_time_compute, muennighoff2025s1, yao2023tree_of_thoughts].
@@
In parallel, practical deployments increasingly rely on tool-augmented LLMs: systems that retrieve evidence, query structured databases, or execute programs as part of producing an answer. Retrieval-Augmented Generation (RAG) combines parametric knowledge with retrieved documents to improve factuality and coverage on knowledge-intensive tasks [lewis2020rag]. Agentic methods such as ReAct explicitly interleave reasoning steps with tool actions, enabling models to decompose tasks and gather missing information during inference [yao2022react]. Other approaches train models to decide when and how to call external tools [schick2023toolformer] or to offload computation to symbolic programs [gao2022pal, chen2022program_of_thoughts]. While these methods often improve correctness and traceability, they also introduce important engineering and product trade-offs: more tool calls and longer reasoning traces typically increase latency and cost, and smaller models may struggle with the control logic needed for multi-step tool usage.
@@
Figure 1: LLM Reasoning Evaluation Workflow
@@
Motivated by these trends, this paper presents an LLM Thinking Benchmark focused on evaluating reasoning-and-tool-use effectiveness under realistic constraints. Rather than benchmarking on purely synthetic puzzles, we evaluate two real-world use cases that frequently arise in enterprise and consumer settings: (1) event-centric question answering over graph-structured knowledge and interconnected information sources, and (2) argument understanding and persuasive response generation grounded in open-domain discussions. For the first use case, we draw questions from the Event-QA dataset [costa2020eventqa] and evaluate the ability of an LLM system to interpret questions that naturally map to graph queries and structured lookups. For the second use case, we use the ChangeMyView (CMV) persuasion setting derived from Reddit discussions, which has become a standard benchmark for studying persuasive interaction and argumentation dynamics [tan2016winning_arguments].
@@
To operationalize multi-step reasoning and tool use, we implement a three-stage CoT-style state machine (planning, execution with tool calls, and re-planning/answering) that is conceptually aligned with agentic reasoning frameworks [yao2022react] and test-time deliberation methods [snell2024test_time_compute, muennighoff2025s1]. For the Event-QA setting, we provide tools for querying and exploring a large public knowledge graph (DBpedia) [lehmann2015dbpedia] and for searching and interpreting background text (e.g., Wikipedia). For CMV, we provide a targeted web search tool to retrieve relevant background knowledge for policy and political topics. We then compare (i) a baseline one-shot approach with no planning or tools against (ii) multi-stage planning approaches that invoke tools during inference.
@@
  • We benchmark one-shot prompting versus multi-stage tool-augmented reasoning on two real-world datasets (Event-QA and CMV), capturing both structured knowledge access and persuasive argumentation [costa2020eventqa, tan2016winning_arguments].
@@ II-A Reasoning and inference-time compute
Transformer-based LLMs [vaswani2017attention] have demonstrated emergent reasoning capabilities as scale increases [wang2024historyllm]. Chain-of-Thought prompting [wei2022chain_of_thought] provides a simple mechanism to elicit intermediate reasoning steps, often improving multi-step accuracy. However, the quality of a single generated reasoning trace can be unstable; self-consistency decoding improves robustness by sampling diverse reasoning paths and selecting the most consistent final answer [wang2022self_consistency]. More generally, a growing body of work suggests that increasing inference-time compute—via additional sampling, search, or deliberation—can offer substantial gains, sometimes rivaling parameter scaling [snell2024test_time_compute, muennighoff2025s1]. Tree-of-Thoughts extends CoT by exploring a search tree over partial solutions with self-evaluation, enabling backtracking and lookahead [yao2023tree_of_thoughts]. These lines of work motivate our focus on comparing one-shot answering with multi-stage plan-and-execute pipelines that explicitly allocate additional computation at test time. Additional perspectives on inference-time scaling and recent reasoning-focused models include practitioner surveys and contemporary reasoning-model studies [raschka2025state, deepseek2025r1, li2025ttpo, wang2025underthinking, pan2025coat].
@@ II-B Tool-augmented and program-aided LLMs
A complementary trend is to enhance LLMs with external tools to mitigate limitations in parametric knowledge and to improve grounding. Retrieval-Augmented Generation (RAG) retrieves supporting evidence and conditions generation on it, improving performance on knowledge-intensive NLP tasks [lewis2020rag]. ReAct shows that interleaving reasoning traces with concrete actions (e.g., search) improves task completion and interpretability [yao2022react]. Toolformer proposes self-supervised methods to teach language models when to call APIs and how to integrate their outputs [schick2023toolformer]. Program-aided approaches offload computation to symbolic interpreters: PAL uses program synthesis to solve reasoning problems [gao2022pal], and Program-of-Thoughts prompting separates numerical computation from natural language reasoning [chen2022program_of_thoughts]. Our benchmark is aligned with these approaches in that it evaluates not only final answer quality but also the practical ability of different models to use tools effectively under multi-step control logic, where smaller models may fail due to planning or tool-invocation errors. Several recent benchmarks evaluate LLMs as tool-using agents in interactive settings, including AgentBench and ToolBench, and realistic web environments such as WebArena. These works complement our focus by providing broad environment coverage, while our benchmark emphasizes accuracy–latency–cost trade-offs in two targeted real-world tasks.
@@ II-C Benchmarks for knowledge graphs and persuasion
For structured knowledge access, question answering over knowledge graphs has long been studied, often requiring query construction and entity linking to answer compositional questions. Event-QA targets event-centric question answering and provides query-verbalization pairs that can stress structured retrieval and reasoning over event-focused graphs [costa2020eventqa]. In our work, DBpedia serves as a widely used public knowledge base and SPARQL-accessible graph for structured lookups and graph-style querying [lehmann2015dbpedia]. This setup allows us to test the extent to which tool-assisted reasoning helps models translate natural language questions into structured retrieval steps.
For persuasive argumentation, the ChangeMyView community on Reddit provides a natural environment where users attempt to change an opinion holder’s mind and where success is explicitly signaled. Tan et al. introduced and analyzed this setting, highlighting interaction and linguistic factors correlated with persuasion [tan2016winning_arguments]. CMV has since been widely used for tasks such as persuasion prediction, argument quality modeling, and evidence-grounded counterargument generation. Our benchmark leverages CMV-style prompts to evaluate whether additional tool-based planning and background retrieval improve persuasive response correctness compared to direct one-shot generation.
@@ II-D Cost- and latency-aware evaluation
Finally, practical system choices are increasingly shaped by the interplay between model size, inference-time computation, and tooling overhead. Recent work on test-time scaling emphasizes that allocating more compute at inference (through sampling, search, or deliberation) can be a competitive alternative to parameter scaling [snell2024test_time_compute, muennighoff2025s1]. In tool-augmented systems, each additional tool call adds latency and introduces opportunities for compounding errors. Our experiments explicitly report both accuracy and latency under different planning and tool configurations, enabling cost-aware comparisons between a larger model and a smaller, cheaper alternative.
@@ III Methodology
We implemented the benchmark in Python using LangChain for tool/model integration and LangGraph for deterministic orchestration of multi-step agent workflows [langchain_overview, langgraph_overview, langgraph_graph_api]. We evaluated two real-world settings: (i) event-centric question answering over a knowledge graph (Event-QA) and (ii) persuasive response generation in the ChangeMyView (CMV) setting. For each dataset, we compared a one-shot baseline against a tool-augmented, multi-stage plan–execute–replan agent.
@@ III-A1 Event-QA approaches
    DBpedia: 3-stage planning with DBpedia tools (entity lookup, schema exploration, and SPARQL execution) [lehmann2015dbpedia, dbpedia_sparql_endpoint, w3c_sparql11_query, dbpedia_lookup_github].
III-A2 ChangeMyView approaches
@@
    PlanningSearch: 3-stage planning with topical web search using Tavily Search. We structured retrieval across 10 topic buckets (Politics & Civic Process, Policy Research & Demographics, Economics & Labor, Justice & Law, Foreign Policy & Security, Health & Science, Technology & Standards, Environment & Energy, Immigration & Civil Rights, Housing & Urban Development) [tavily_search_endpoint].
@@ III-B LangGraph approaches
We implemented two LangGraph controller patterns, shown in Fig. 2. The left pipeline is the one-shot baseline, and the right pipeline is the multi-stage planning agent [langgraph_overview, langgraph_graph_api].
Figure 2: The two LangGraph approaches evaluated. Left: a one-shot baseline where the LLM answers directly. Right: a plan–execute–replan pipeline where the LLM plans, invokes tools during execution, and then answers or revises the plan.
@@ III-C LangGraph tools
    Web search tool: returns ranked snippets from search results. For Event-QA we configured it for Wikipedia-focused retrieval; for CMV we used Tavily Search for topical background retrieval [tavily_search_endpoint].
@@
    DBpedia SPARQL query tool: executes SPARQL 1.1 queries against the DBpedia endpoint [dbpedia_sparql_endpoint, w3c_sparql11_query].
@@
    DBpedia resource/entity lookup tool: resolves surface forms (e.g., entity names) into DBpedia URIs for use in SPARQL queries [dbpedia_lookup_github].
@@
    DBpedia schema explorer tool: retrieves relevant ontology/types/properties for an entity to guide query construction (e.g., which predicates support filtering/counting) [lehmann2015dbpedia].
@@ III-D LLMs used
We compared two OpenAI models: GPT-4o and GPT-4o-mini [openai_gpt4o_model, openai_gpt4o_mini_model]. Table I summarizes their documented context limits and token pricing. Token prices can vary by service tier (e.g., batch vs. standard vs. priority); we cite the official pricing documentation for the values reported [openai_api_pricing]. Because OpenAI does not publish official parameter counts for these models, we report commonly cited third-party estimates [claudeArtifact0ecdfb83, howarth2025_gptParameters].
Table I reports OpenAI’s published per-token prices for the Standard processing tier at the time of access.[openai_api_pricing].
@@ IV-A Data
We evaluated on two labeled datasets from prior work: (i) Event-QA for event-centric question answering over knowledge graphs [costa2020eventqa], and (ii) ChangeMyView (CMV) for persuasion/argumentation in good-faith online discussions [tan2016winning_arguments]. Due to the high runtime cost of multi-stage tool-augmented inference (Section III), we conducted experiments on fixed-size subsets.
@@ IV-B Evaluation metrics
IV-B1 ChangeMyView (CMV) automatic evaluation
Because CMV responses are open-ended, we operationalize correctness using ROUGE-1 F-measure (unigram overlap) against the set of human reference responses [lin2004rouge]. For each example $i$, we compute ROUGE-1 F between the generated response $\hat{y}_{i}$ and each reference $r\in\mathcal{R}_{i}$, take the best match
@@ IV-C1 EventQA results
Figure 3: EventQA - GPT-4o
Figure 4: EventQA - GPT-4o-mini
@@
Figure 5: EventQA - Best Overall
TABLE II: EventQA: Comparison of Approach and Model Performance, Speed, and Optimal Configuration

### v2

# When Do Tools and Planning Help Large Language Models Think? A Cost- and Latency-Aware Benchmark

@@ Abstract
Modern large language models (LLMs) increasingly rely on inference-time planning and external tools to improve reasoning. We benchmark this behavior on two real-world settings: event-centric question answering over graph-structured knowledge (Event-QA) and persuasive response generation in Reddit ChangeMyView (CMV). Using LangChain and LangGraph, we compare a one-shot baseline against a plan–execute–replan agent equipped with task-specific tools (DBpedia SPARQL Protocol and RDF Query Language (SPARQL)/lookup/schema exploration, Wikipedia-focused retrieval, and topical web search). We evaluate on 60 examples each from Event-QA and CMV (3 splits of 20), and report both mean end-to-end latency and per-example token costs. We evaluate GPT-4o and GPT-4o-mini under identical workflows and report accuracy and end-to-end latency. On Event-QA, the best tool-augmented configuration improves accuracy (e.g., 47.5% $\rightarrow$ 67.5% for GPT‑4o) while increasing latency by orders of magnitude ($\sim$8s $\rightarrow$ $\sim$317s per example). On CMV, one-shot prompting is strongest (e.g., GPT‑4o‑mini achieves 75% accuracy at $\sim$6s), and planning+search increases latency substantially without consistent gains. However, complex multi-tool orchestration exposes failure modes where the smaller model degrades. Overall, the findings highlight the need for task-specific, cost-aware choices of both model size and agent/tooling complexity.
@@ I Introduction
Large Language Models (LLMs) based on Transformer architectures have rapidly evolved into general-purpose systems for text understanding, generation, and reasoning [25, 29]. Beyond surface-level language fluency, recent development has emphasized the ability of LLMs to perform multi-step reasoning (planning intermediate steps and integrating external evidence)—i.e., to perform multi-step reasoning, plan intermediate steps, and reliably integrate external information sources when answering complex real-world questions. A common and influential approach is Chain-of-Thought (CoT) prompting, which elicits intermediate reasoning traces that can substantially improve performance on multi-step problems [31]. Subsequent work has shown that inference-time strategies such as sampling multiple reasoning paths and aggregating consistent answers (self-consistency) can further improve reliability without changing model parameters [27]. More broadly, recent lines of research argue that test-time compute—allocating additional deliberation or search at inference—can sometimes yield improvements comparable to or greater than scaling model size [22, 15, 32].
@@
In parallel, practical deployments increasingly rely on tool-augmented LLMs: systems that retrieve evidence, query structured databases, or execute programs as part of producing an answer. Retrieval-Augmented Generation (RAG) combines parametric knowledge with retrieved documents to improve factuality and coverage on knowledge-intensive tasks [12]. Agentic methods such as ReAct explicitly interleave reasoning steps with tool actions, enabling models to decompose tasks and gather missing information during inference [33]. Other approaches train models to decide when and how to call external tools [21] or to offload computation to symbolic programs [6, 1]. While these methods often improve correctness and traceability, they also introduce important engineering and product trade-offs: more tool calls and longer reasoning traces typically increase latency and cost, and smaller models may struggle with the control logic needed for multi-step tool usage.
@@
Figure 1: LLM Reasoning Evaluation Workflow, detailing the process of planning, tool integration, and execution.
@@
Motivated by these trends, this paper presents an LLM Thinking Benchmark focused on evaluating reasoning-and-tool-use effectiveness under realistic constraints. Rather than benchmarking on purely synthetic puzzles, we evaluate two real-world use cases that frequently arise in enterprise and consumer settings: (1) event-centric question answering over graph-structured knowledge and interconnected information sources, and (2) argument understanding and persuasive response generation grounded in open-domain discussions. For the first use case, we draw questions from the Event-QA dataset [2] and evaluate the ability of an LLM system to interpret questions that naturally map to graph queries and structured lookups. For the second use case, we use the ChangeMyView (CMV) persuasion setting derived from Reddit discussions, which has become a standard benchmark for studying persuasive interaction and argumentation dynamics [23].
@@
To operationalize multi-step reasoning and tool use, we implement a three-stage CoT-style state machine (planning, execution with tool calls, and re-planning/answering) that is conceptually aligned with agentic reasoning frameworks [33] and test-time deliberation methods [22, 15]. For the Event-QA setting, we provide tools for querying and exploring a large public knowledge graph (DBpedia) [11] and for searching and interpreting background text (e.g., Wikipedia). For CMV, we provide a targeted web search tool to retrieve relevant background knowledge for policy and political topics. We then compare (i) a baseline one-shot approach with no planning or tools against (ii) multi-stage planning approaches that invoke tools during inference.
@@
  • We benchmark one-shot prompting versus multi-stage tool-augmented reasoning on two real-world datasets (Event-QA and CMV), capturing both structured knowledge access and persuasive argumentation [2, 23].
@@ II-A Reasoning and Inference-Time Compute
Transformer-based LLMs [25] have demonstrated emergent reasoning capabilities as scale increases [29]. Chain-of-Thought prompting [31] provides a simple mechanism to elicit intermediate reasoning steps, often improving multi-step accuracy. However, the quality of a single generated reasoning trace can be unstable; self-consistency decoding improves robustness by sampling diverse reasoning paths and selecting the most consistent final answer [27]. More generally, a growing body of work suggests that increasing inference-time compute—via additional sampling, search, or deliberation—can offer substantial gains, sometimes rivaling parameter scaling [22, 15]. Tree-of-Thoughts extends CoT by exploring a search tree over partial solutions with self-evaluation, enabling backtracking and lookahead [32]. These lines of work motivate our focus on comparing one-shot answering with multi-stage plan-and-execute pipelines that explicitly allocate additional computation at test time. Additional perspectives on inference-time scaling and recent reasoning-focused models include practitioner surveys and contemporary reasoning-model studies [20, 5, 13, 28, 19].
@@ II-B Tool-augmented and Program-Aided LLMs
A complementary trend is to enhance LLMs with external tools to mitigate limitations in parametric knowledge and to improve grounding. Retrieval-Augmented Generation (RAG) retrieves supporting evidence and conditions generation on it, improving performance on knowledge-intensive NLP tasks [12]. ReAct shows that interleaving reasoning traces with concrete actions (e.g., search) improves task completion and interpretability [33]. Toolformer proposes self-supervised methods to teach language models when to call APIs and how to integrate their outputs [21]. Program-aided approaches offload computation to symbolic interpreters: PAL uses program synthesis to solve reasoning problems [6], and Program-of-Thoughts prompting separates numerical computation from natural language reasoning [1]. Our benchmark is aligned with these approaches in that it evaluates not only final answer quality but also the practical ability of different models to use tools effectively under multi-step control logic, where smaller models may fail due to planning or tool-invocation errors. Several recent benchmarks evaluate LLMs as tool-using agents in interactive settings, including AgentBench and ToolBench, and realistic web environments such as WebArena. These works complement our focus by providing broad environment coverage, while our benchmark emphasizes accuracy–latency–cost trade-offs in two targeted real-world tasks.
@@ II-C Benchmarks for Knowledge Graphs and Persuasion
For structured knowledge access, question answering over knowledge graphs has long been studied, often requiring query construction and entity linking to answer compositional questions. Event-QA targets event-centric question answering and provides query-verbalization pairs that can stress structured retrieval and reasoning over event-focused graphs [2]. In our work, DBpedia serves as a widely used public knowledge base and SPARQL-accessible graph for structured lookups and graph-style querying [11]. This setup allows us to test the extent to which tool-assisted reasoning helps models translate natural language questions into structured retrieval steps.
For persuasive argumentation, the CMV community on Reddit provides a natural environment where users attempt to change an opinion holder’s mind and where success is explicitly signaled. Tan et al. introduced and analyzed this setting, highlighting interaction and linguistic factors correlated with persuasion [23]. CMV has since been widely used for tasks such as persuasion prediction, argument quality modeling, and evidence-grounded counterargument generation. Our benchmark leverages CMV-style prompts to evaluate whether additional tool-based planning and background retrieval improve persuasive response correctness compared to direct one-shot generation.
@@ II-D Cost- and Latency-Aware Evaluation
Finally, practical system choices are increasingly shaped by the interplay between model size, inference-time computation, and tooling overhead. Recent work on test-time scaling emphasizes that allocating more compute at inference (through sampling, search, or deliberation) can be a competitive alternative to parameter scaling [22, 15]. In tool-augmented systems, each additional tool call adds latency and introduces opportunities for compounding errors. Our experiments explicitly report both accuracy and latency under different planning and tool configurations, enabling cost-aware comparisons between a larger model and a smaller, cheaper alternative.
@@ III Methodology
We implemented the benchmark in Python using LangChain for tool/model integration and LangGraph for deterministic orchestration of multi-step agent workflows [9, 10, 8]. We evaluated two real-world settings: (i) event-centric question answering over a knowledge graph (Event-QA) and (ii) persuasive response generation in the ChangeMyView (CMV) setting. For each dataset, we compared a one-shot baseline against a tool-augmented, multi-stage plan–execute–replan agent.
@@ III-A1 Event-QA Approaches
    DBpedia: 3-stage planning with DBpedia tools (entity lookup, schema exploration, and SPARQL execution) [11, 3, 26, 4].
III-A2 CMV Approaches
@@
    PlanningSearch: 3-stage planning with topical web search using Tavily Search. We structured retrieval across 10 topic buckets (Politics & Civic Process, Policy Research & Demographics, Economics & Labor, Justice & Law, Foreign Policy & Security, Health & Science, Technology & Standards, Environment & Energy, Immigration & Civil Rights, Housing & Urban Development) [24].
@@ III-B LangGraph Approaches
We implemented two LangGraph controller patterns, shown in Fig. 2. The left pipeline is the one-shot baseline, and the right pipeline is the multi-stage planning agent [10, 8].
Figure 2: The two LangGraph approaches evaluated. Left: a one-shot baseline where the LLM answers directly. Right: a plan–execute–replan pipeline where the LLM plans, invokes tools during execution, and then answers or revises the plan (e.g., invokes tools for knowledge retrieval or execution).
@@ III-C LangGraph Tools
    Web search tool: returns ranked snippets from search results. For Event-QA we configured it for Wikipedia-focused retrieval; for CMV we used Tavily Search for topical background retrieval [24].
@@
    DBpedia SPARQL query tool: executes SPARQL 1.1 Protocol and RDF Queries against the DBpedia endpoint [3, 26].
@@
    DBpedia resource/entity lookup tool: resolves surface forms (e.g., entity names) into DBpedia URIs for use in SPARQL queries [4].
@@
    DBpedia schema explorer tool: retrieves relevant ontology/types/properties for an entity to guide query construction (e.g., which predicates support filtering/counting) [11].
@@ III-D LLMs Used
We compared two OpenAI models: GPT-4o and GPT-4o-mini [17, 16]. Table I summarizes their documented context limits and token pricing. Token prices can vary by service tier (e.g., batch vs. standard vs. priority); we cite the official pricing documentation for the values reported [18]. Because OpenAI does not publish official parameter counts for these models, we report commonly cited third-party estimates [30, 7].
Table I reports OpenAI’s published per-token prices for the Standard processing tier at the time of access.[18].
@@ IV-A Data
We evaluated on two labeled datasets from prior work: (i) Event-QA for event-centric question answering over knowledge graphs [2], and (ii) CMV for persuasion/argumentation in good-faith online discussions [23]. Due to the high runtime cost of multi-stage tool-augmented inference (Section III), we conducted experiments on fixed-size subsets.
@@ IV-B Evaluation Metrics
IV-B1 CMV Automatic Evaluation
Because CMV responses are open-ended, we operationalize correctness using ROUGE-1 F-measure (unigram overlap) against the set of human reference responses [14]. For each example $i$, we compute ROUGE-1 F between the generated response $\hat{y}_{i}$ and each reference $r\in\mathcal{R}_{i}$, take the best match
@@ IV-C1 Event-QA Results
Figure 3: Event-QA - GPT-4o
Figure 4: Event-QA - GPT-4o-mini
@@
Figure 5: Event-QA - Mixed Results - Planning with DBPedia is most acurate (GPT-4o, 67.5% Accuracy at 317s latency) while Planning with Wikipedia Search has balanced accuracy vs speed (GPT-4o-mini, 55% Accuracy at 84s latency)
TABLE II: Event-QA: Comparison of Approach and Model Performance, Speed, and Optimal Configuration. Compares the best overall accuracy and latency for the optimal configuration per model against the NoPlanning baseline.